a chemist needs 110 milliliters of an 80% solution but has only 76% and 98% solutions available.

Answer by addingup(3677)

Let the amount of solution at 76% be x. Then, the amount of solution at 98% will be 110-x since 110 is the total quantity desired.
0.76x+0.98(110-x)= 0.80(110)
0.76x+107.8-0.98x = 88
-0.22x = -19.8 Divide both sides by -0.22 and remember that -/- = +
x = 90
You need 90ml of 76% and 110-90 = 20 ml of the 98%